A body of mass 4 kg, projected at an angle of 65° from the ground with an initial velocity of 36 m/s, acceleration due to gravity is g = 9.81 m/s2, determine:

a) time of flight 


b) maximum horizontal distance covered 


c) maximum vertical distance covered 



1
Expert's answer
2022-05-22T14:38:11-0400

"t=\frac{2v\sin\alpha}g=6.65~s,"

"s=\frac{v\sin 2\alpha}g=101.2~m,"

"h=\frac{v^2\sin^2\alpha}{2g}=54.3~m."